There is a probability of a man to live 20 more years and his wife to live 20 more years.
What is the probability of;
1) both to live 20 more years?
2) At least one person to live 20 more years?
3) Only the man to live 20 more years?
4) Only his wife to live 20 more years?
5) man or the wife to live 20 more years?
6) At least one person NOT to live 20 more years

Answer by ikleyn(52781)

.
Your formulation of the given part is incorrect (which demonstrates that you do not understand what you are posting for . . . (unfortunately) )

The correct formulation of the given part is THIS:

    The probability of a man to live 20 more years is X and the probability of his wife to live 20 more years is Y.


The partial answers are:

1)  1 - (1-X)*(1-Y) = X + Y - XY.


2)  X*(1-Y) + Y*(1-X) = X + Y - 2*XY.


3)  X*(1-Y).

------------------
This problem was posted to the forum, probably, one year ago.

See the link
https://www.algebra.com/algebra/homework/Probability-and-statistics/Probability-and-statistics.faq.question.1105633.html

https://www.algebra.com/algebra/homework/Probability-and-statistics/Probability-and-statistics.faq.question.1105633.html


and my solution there.